# LoKI: Low-damage Knowledge Implanting of Large Language Models

### v1

@@ A.1.1 Step setting
We examined the effect of the Riemann approximation step $m$ on the difference of trainable node positions determined by the Layer-Balanced Strategy ($q{=}10$). In particular, we compare the outcomes using $m{=}7$ and $m{=}20$ on Qwen2.5-0.5B-Instruct. Table 7 reports the similarity between the resulting trainable node positions under these two configurations.
@@ A.1.2 Number of samples
Table 8: The similarity of each layer and the total similarity of trainable nodes obtained using two different MMLU sample selection methods on Qwen2.5-0.5B-Instruction.
We also investigated the impact of using different numbers of MMLU samples in the KVA process on the final trainable nodes. Specifically, we compared the similarity between using only the first 50 samples for each MMLU subset and using all MMLU samples. Note that we compare at $m{=}7$ and $q{=}10$. Following the definition of similarity in the previous section, we report the node similarity obtained by two methods on Qwen2.5-0.5B-Instruct and Llama3.1-8B-Instruct in Tables 8 and 9, respectively.
@@ A.2 Discussions
As mentioned in Section 3.1, the KVA process involves a one-time computational overhead, which can be considered from both the perspectives of time and computational cost. Here, we discuss potential directions for optimizing both types of overhead.
@@ Appendix B KVA Heatmaps
B.1 Qwen2.5-0.5B-Instruct
Figure 3: Heatmaps of top $10\%$ KVA results across all 24 FFN layers of Qwen2.5-0.5B-Instruct. The upper (red-tinted) map highlights the distribution of high-contribution node positions, while the lower (blue-tinted) map highlights the distribution of low-contribution node positions. Color intensity (log-scale) indicates the density of neurons in each category, with darker colors representing higher density.
B.2 Llama2-7B
Figure 4: Heatmaps of top $5\%$ KVA results across all 32 FFN layers of Llama2-7B, following the same visualization settings as in Fig. 3.
@@ Appendix C Implementation of LoKI Linear
We provide the PyTorch implementation code for LoKI Linear, and readers can access the complete project code on our public GitHub repository. Based on this implementation, we performed all the LoKI-based model training in the paper. Our code has been validated for use with Llama-Factory [^2] URL omitted for double-blind review.
and supports integration with LoRA.
{minted}
[ frame=lines, linenos, breaklines, fontsize=]python import torch import torch.nn as nn
class LoKILinear(nn.Module): def __init__(self, original_linear, target_pos): super().__init__() self.out_features = original_linear.out_features self.in_features = original_linear.in_features self.active_pos = sorted(target_pos) self.frozen_pos = [ i for i in range(self.out_features) if i not in self.active_pos ]
# Parameter validation if not all(0 <= idx < self.out_features for idx in self.active_pos): raise ValueError(f"Activation indices must be within [0, self.out_features - 1]") if len(self.active_pos) != len(set(self.active_pos)): raise ValueError("Activation indices contain duplicate values") self.active = nn.Linear(self.in_features, len(self.active_pos), bias=False) self.frozen = nn.Linear(self.in_features, len(self.frozen_pos), bias=False) # Split the weight matrix W = original_linear.weight.data self.active.weight = nn.Parameter(W[self.active_pos].clone(), requires_grad=True) self.frozen.weight = nn.Parameter(W[self.frozen_pos].clone(), requires_grad=False)
# Handle bias if original_linear.bias is not None: b = original_linear.bias.data self.active_bias = nn.Parameter( b[self.active_pos].clone(), requires_grad=True ) self.frozen_bias = nn.Parameter( b[self.frozen_pos].clone(), requires_grad=False ) else: self.register_parameter("active_bias", None) self.register_parameter("frozen_bias", None)
# Pre-generate index mapping index_map = torch.empty(self.out_features, dtype=torch.long) index_map[self.active_pos] = torch.arange(len(self.active_pos)) index_map[self.frozen_pos] = torch.arange(len(self.frozen_pos)) + len( self.active_pos ) self.register_buffer("index_map", index_map)
def forward(self, x): active_out = self.active(x) # Compute active part via submodule frozen_out = self.frozen(x) # Fixed part output = torch.cat([active_out, frozen_out], dim=-1)
# Add combined bias if self.active_bias is not None: bias = torch.cat([self.active_bias, self.frozen_bias], dim=0) output += bias.unsqueeze(0).unsqueeze(0) # Broadcast bias to all batches and sequence positions
# Reorder output using pre-generated indices return output.gather( dim=-1, index=self.index_map.view(1, 1, -1).expand( output.size(0), output.size(1), -1 ), )
@@ Appendix D Experiment Setups for the LB Reranker Dataset
Model Learning Rate Batch Size Epochs LR Scheduler WarmUp Ratio
LoKI($q{=}5$) 1.0e-5 1 1 cosine 0.01
LoKI($q{=}10$) 1.0e-5
LoKI($q{=}20$) 5.0e-6
LoKI($q{=}30$) 4.0e-6
Table 10: Training hyperparameters on LB Reranker Dataset.
@@
We use the BEIR evaluation code [^3] URL omitted for double-blind review. provided by the dataset provider to assess our models. Specifically, we evaluate on 9 datasets from BEIR: Arguana, Dbpedia-entity, FiQA, NFCorpus, SCIDOCS, SciFact, TREC-COVID-v2, ViHealthQA, and Webis-Touche2020. For each dataset, we evaluate on a subset of the queries (the first 250).
@@ Appendix E Experiment Setups for the ToolACE Function-Calling Dataset
Model Learning Rate Batch Size Epochs LR Scheduler WarmUp Ratio
LoKI($q{=}10$) 9.0e-06 4 3 cosine 0.1
LoKI($q{=}20$)
LoKI($q{=}30$)
Table 11: Training hyperparameters on ToolACE Function-Calling Dataset.
@@
LoKI*($q{=}30$) 5.0*e-04 32 64
Table 12: Training hyperparameters on ToolACE Function-Calling Dataset when using LoRA. The parameters not mentioned in the table will continue to use the parameters in Table 11.
@@
As mentioned in the main text, we use the publicly available evaluation codebase for the Berkeley Function Calling Leaderboard (BFCL) [^4] URL omitted for double-blind review..
@@ Appendix F Benchmark Execution Details
To ensure a fair comparison of model performance, we use OpenCompass [^5] URL omitted for double-blind review. to evaluate all the models discussed in this paper, including the two base models: Llama3.1-8B-Instruct and Qwen2.5-0.5B-Instruction. The version of the benchmarks can be viewed in Table 13. Specifically, for IFEval, we calculate the average of four indicators. The scores for the remaining benchmarks are reported directly.
@@
Appendix G Performance on the BEIR Benchmark for G-H/L
Table 14: Comparison with the full parameter fine-tuning model on BEIR.
@@ Appendix G Performance on the BEIR Benchmark for G-H/L
G-H($q{=}10$) -1.0 -2.8 -1.0 -1.6 0.4 -1.2 0.0 -1.0
G-L($q{=}10$) -4.5 -4.8 -4.5 -2.8 -2.6 -3.0 -2.6 -3.0
Table shows the percentage difference in performance compared to the full parameter fine-tuning model across standard retrieval metrics.
@@ Appendix H Limitations and Future Works
[…]
Appendix I Broader Impacts
Positive Societal Impacts. LoKI enables LLMs to be adapted to new tasks with minimal loss of pre-trained general knowledge. By preserving general capabilities, LoKI promotes the development of multi-domain models that retain versatility across a wide range of tasks, potentially reducing the need to maintain and manage a large number of specialized models. This can foster more unified and maintainable AI systems with consistent behavior across applications.
Negative Societal Impacts and Risks. As with any technique that enhances the efficiency and accessibility of LLM customization, LoKI may inadvertently lower the barrier for misuse. Specifically, malicious actors could leverage LoKI to implant biased, harmful, or deceptive capabilities into powerful language models while preserving general fluency, potentially making such manipulations harder to detect. Additionally, while LoKI protects general knowledge, the process of knowledge implantation could still result in subtle shifts in behavior that are difficult to audit.
Moreover, the reliance on pretrained LLMs—often trained on opaque or uncurated datasets—means that any biases present in the original model are likely to persist or be reinforced during adaptation. LoKI does not directly address these foundational concerns, and care must be taken to ensure its application aligns with ethical and fairness principles.
Mitigation Strategies. To mitigate risks, we encourage the use of LoKI in conjunction with auditing tools, attribution techniques, and interpretability methods to ensure transparency in model modifications. Furthermore, deployment should be accompanied by rigorous evaluation on fairness, robustness, and safety metrics, especially when the adapted models are intended for high-stakes applications.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: In the abstract and introduction we introduce the current situation of PEFT and CF. We claim that our work provides a PEFT approach to alleviate CF while maintaining downstream task performance.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss the limitations of our work in Appendix H.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We include theoretical justifications for our method in Section 3.1.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We explain our experiments in Section 4 and provide the complete set of hyper-parameters in Appendix D and Appendix E.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We provide a zip file containing the code for our experiments, and the GitHub repository will be made public.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so "No" is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We provide information about the training details in Appendix D and Appendix E.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: We do not provide error bars, but we explained the hyperparameter selection criteria for all experiments in Appendix D and Appendix E. We only adjusted the learning rate to suit various quotas and documented the best results.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: All these information can be found in Appendix D and Appendix E.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research conducted in this paper conforms, in every respect, with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We discuss broader impacts in Appendix J.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: We do not release any models.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We work with open-source models and datasets that are publicly available, and we cited them properly.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We provide code and instructions on how to run the code to reproduce our results.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: The method proposed in this paper mainly focuses on LLMs as the execution object.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v2

@@ A.1 Exploration of Different Settings
We examined the effect of the Riemann approximation step $m$ on the difference of trainable node positions determined by the Layer-Balanced Strategy (q=10q{=}10). In particular, we compare the outcomes using m=7m{=}7 and m=20m{=}20 on Qwen2.5-0.5B-Instruct. Table 6 reports the similarity between the resulting trainable node positions under these two configurations.
@@
Table 7: The similarity of each layer and the total similarity of trainable nodes obtained using two different MMLU sample selection methods on Qwen2.5-0.5B-Instruct.
We also investigated the impact of using different numbers of MMLU samples in the KVA process on the final trainable nodes. Specifically, we compared the similarity between using only the first 50 samples for each MMLU subset and using all MMLU samples. Note that we compare at m=7m{=}7 and q=10q{=}10. Following the definition of similarity in the previous section, we report the node similarity obtained by two methods on Qwen2.5-0.5B-Instruct and Llama3.1-8B-Instruct in Tables 7 and 8, respectively.
@@ A.2 Discussions
As mentioned in Section Analysing, the KVA process involves a one-time computational overhead, which can be considered from both the perspectives of time and computational cost. Here, we discuss potential directions for optimizing both types of overhead.
@@ Appendix B KVA Heatmaps
Figure 3: Heatmaps of top $10\%$ KVA results across all 24 layers of Qwen2.5-0.5B-Instruct. The upper (red-tinted) map highlights the distribution of high-contribution node positions, while the lower (blue-tinted) map highlights the distribution of low-contribution node positions. Color intensity (log-scale) indicates the density of neurons in each category, with darker colors representing higher density.
Figure 4: Heatmaps of top $5\%$ KVA results across all 32 layers of Llama2-7B, following the same visualization settings as in Fig. 3.
Figure 5: Heatmaps of top $5\%$ KVA results across all 32 layers of Llama3.1-8B-Instruct, following the same visualization settings as in Fig. 3.
@@ Appendix C Implementation of LoKI Linear
We provide the PyTorch implementation code for LoKI Linear, and readers can access the complete project code on our public GitHub repository. Based on this implementation, we performed all the LoKI-based model training in the paper. Our code has been validated for use with Llama-Factory and supports integration with LoRA.
Listing 1: PyTorch implementation of LoKI Linear
1import torch
2import torch.nn as $n$
3
4
5class LoKILinear($n$.Module):
6 def __init__(self, original_linear, target_pos):
7 super().__init__()
8 self.out_features = original_linear.out_features
9 self.in_features = original_linear.in_features
10 self.active_pos = sorted(target_pos)
11 self.frozen_pos = [
12 i for i in range(self.out_features) if i not in self.active_pos
13 ]
14
15 # Parameter validation
16 if not all(0 <= idx < self.out_features for idx in self.active_pos):
17 raise ValueError(f"Activation indices must be within [0, {self.out_features - 1}]")
18 if len(self.active_pos) != len(set(self.active_pos)):
19 raise ValueError("Activation indices contain duplicate values")
20 self.active = $n$.Linear(self.in_features, len(self.active_pos), bias=False)
21 self.frozen = $n$.Linear(self.in_features, len(self.frozen_pos), bias=False)
22 # Split the weight matrix
23 W = original_linear.weight.data
24 self.active.weight = $n$.Parameter(W[self.active_pos].clone(), requires_grad=True)
25 self.frozen.weight = $n$.Parameter(W[self.frozen_pos].clone(), requires_grad=False)
26
27 # Handle bias
28 if original_linear.bias is not None:
29 b = original_linear.bias.data
30 self.active_bias = $n$.Parameter(
31 b[self.active_pos].clone(), requires_grad=True
32 )
33 self.frozen_bias = $n$.Parameter(
34 b[self.frozen_pos].clone(), requires_grad=False
35 )
36 else:
37 self.register_parameter("active_bias", None)
38 self.register_parameter("frozen_bias", None)
39
40 # Pre-generate index mapping
41 index_map = torch.empty(self.out_features, dtype=torch.long)
42 index_map[self.active_pos] = torch.arange(len(self.active_pos))
43 index_map[self.frozen_pos] = torch.arange(len(self.frozen_pos)) + len(
44 self.active_pos
45 )
46 self.register_buffer("index_map", index_map)
47
48 def forward(self, x):
49 active_out = self.active(x) # Compute active part via submodule
50 frozen_out = self.frozen(x) # Fixed part
51 output = torch.cat([active_out, frozen_out], dim=-1)
52
53 # Add combined bias
54 if self.active_bias is not None:
55 bias = torch.cat([self.active_bias, self.frozen_bias], dim=0)
56 output += bias.unsqueeze(0).unsqueeze(0) # Broadcast bias to all batches and sequence positions
57
58 # Reorder output using pre-generated indices
59 return output.gather(
60 dim=-1,
61 index=self.index_map.view(1, 1, -1).expand(
62 output.size(0), output.size(1), -1
63 ),
64 )
@@ Appendix D Experiment Setups for the LB Reranker Dataset
Model Learning Rate
Batch
Size
Epochs
LR
Scheduler
WarmUp
Ratio
LoKI(q=5q{=}5) $\times10$-5 1 1 cosine 0.01
LoKI(q=10q{=}10) $\times10$-5
[…]
Table 9: Training hyperparameters on LB Reranker Dataset. LR denotes Learning Rate.
Model Learning Rate LoRA rank LoRA alpha
DoRA $\times10$-5 8 16
PiSSA
CorDA
Table 10: Training hyperparameters on LB Reranker Dataset when using DoRA, PiSSA, and CorDA. The parameters not mentioned in the table use the same parameters as in Table 9.
@@
We use the BEIR evaluation code provided by the dataset provider to assess our models. Specifically, we evaluate on 9 datasets from BEIR: Arguana, Dbpedia-entity, FiQA, NFCorpus, SCIDOCS, SciFact, TREC-COVID-v2, ViHealthQA, and Webis-Touche2020. For each dataset, we evaluate on a subset of the queries (the first 250).
@@ Appendix E Experiment Setups for the ToolACE Function-Calling Dataset
Model Learning Rate
Batch
Size
Epochs
LR
Scheduler
WarmUp
Ratio
[…]
LoKI(q=20q{=}20)
LoKI(q=30q{=}30)
Table 11: Training hyperparameters on ToolACE Function-Calling Dataset. LR denotes Learning Rate.
@@
[…]
Table 12: Training hyperparameters on ToolACE Function-Calling Dataset for LoKI*(q=30q{=}30). The parameters not mentioned in the table use the same parameters as in Table 11.
Model Learning Rate LoRA rank LoRA alpha
DoRA $\times10$-4 16 32
PiSSA
Table 13: Training hyperparameters on ToolACE Function-Calling Dataset when using DoRA and PiSSA. The parameters not mentioned in the table use the same parameters as in Table 11.
@@
As mentioned in the main text, we use the publicly available evaluation codebase for the Berkeley Function Calling Leaderboard (BFCL).
@@ Appendix F Benchmark Execution Details
To ensure a fair comparison of model performance, we use OpenCompass to evaluate all the models discussed in this paper, including the two base models: Llama3.1-8B-Instruct and Qwen2.5-0.5B-Instruction. The version of the benchmarks can be viewed in Table 14. Specifically, for IFEval, we calculate the average of four indicators. The scores for the remaining benchmarks are reported directly.
@@
Appendix G Details for Ablation Studies
Model Trivia QA GSM 8K Hella Swag Wino Grande Human Eval IF Eval
Avg(↓)
LoKI 20.53 37.98 21.53 47.04 24.39 33.39 8.86%
G-H
(q=10q{=}10)
11.04 29.34 10.23 17.68 25.00 26.79 39.04%
G-L
(q=10q{=}10)
9.83 30.93 8.48 43.09 24.39 27.72 30.48%
Table 15: Performance comparison of different methods on general task benchmarks.
Model Trivia QA GSM 8K Hella Swag Wino Grande Human Eval IF Eval
Avg(↓)
Llama 65.77 84.46 73.85 62.98 68.29 79.76 NaN
S-H 62.32 15.16 23.54 60.93 43.29 59.73 36.73%
S-L 62.05 57.85 73.84 60.3 54.88 74.15 11.35%
Table 16: Performance of two suppression strategies on benchmarks when q=1q{=}1.
@@ Appendix G Details for Ablation Studies
G-H
(q=10q{=}10)
-1.0 -2.8 -1.0 -1.6 0.4 -1.2 0.0 -1.0
G-L
(q=10q{=}10)
-4.5 -4.8 -4.5 -2.8 -2.6 -3.0 -2.6 -3.0
Table 17: Comparison with the full parameter fine-tuning model on BEIR. Table shows the percentage difference in performance compared to the full parameter fine-tuning model across standard retrieval metrics.
@@ Appendix H Limitations and Future Works
While our method demonstrates competitive performance across experimental evaluations, several limitations warrant discussion. First, the effectiveness of LoKI is influenced by the hyperparameter $q$, which controls the proportion of trainable nodes. While we provide empirical guidelines (e.g., $q{=}10\text{--}30$), optimal values vary across tasks and models. A more robust, adaptive mechanism for quota allocation would improve LoKI’s general usability. Second, the one-time computational overhead introduced by KVA has significant potential for optimization, especially in implementing code. As shown in Appendix A, there is still a trade-off to be addressed between specific parameter settings and overall performance. Additionally, when integrating LoRA within the LoKI framework, the interaction between LoRA training parameters and the LoKI $q$ value remains underexplored. Finally, the attribution techniques employed in the analysing phase offer further opportunities for enhancement. Future work could focus on improving both the accuracy and efficiency of these attribution methods.